OpenStudy (anonymous):

xy = –16. What is the constant of variation for this relationship? A.-6 B.-1.6 C.1/4 D.-0.16

OpenStudy (anonymous):

and i will give medal to the best answer and explanation

jimthompson5910 (jim_thompson5910):

General form of inverse variation equations x*y = k k is the constant of variation

jimthompson5910 (jim_thompson5910):

x*y = k can be rewritten into \(\LARGE y = \frac{k}{x}\) after solving for x

OpenStudy (anonymous):

ok i got it written down what next?

OpenStudy (anonymous):

can you give me the answer so i can check my work? then i will give you a medal.

OpenStudy (anonymous):

jim_thomsan?

jimthompson5910 (jim_thompson5910):

the answer is practically given to you compare x*y = k to the equation you have x*y = -16 you'll see what the value of k must be

OpenStudy (anonymous):

so -16 is the answer

OpenStudy (anonymous):

no not that

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

jim?

OpenStudy (anonymous):

jim_thopson 5910?

jimthompson5910 (jim_thompson5910):

yeah k = -16

jimthompson5910 (jim_thompson5910):

so -16 is the constant of variation